OpenStudy (anonymous):

WILL GIVE MEDAL PLEASE HELP :) Which sets of numbers are closed under subtraction? Choose all answers that are correct. A. whole numbers B. negative integers C. integers D. rational numbers

OpenStudy (anonymous):

@texaschic101

myininaya (myininaya):

Do you know the definition of when an operation is consider closed?

OpenStudy (anonymous):

no :/

myininaya (myininaya):

Ok an example of a set that closed under addition is the set of integers This means any two integers I take from the set of integers and add them I will still get a number in the set of integers. examples for my example: -1+(-1)=-2<--still an integer -1+0=-1<--still an integer An example of a set not being closed under addition is the set of irrationals. sqrt(2) is irrational but sqrt(2)+(-sqrt(2))=0 and 0 is not irrational

myininaya (myininaya):

So your operation is subtraction do you know what the wholes numbers are?

OpenStudy (anonymous):

yeah

myininaya (myininaya):

can you think of example where the set is not closed under subtraction?

myininaya (myininaya):

If you can think of one example, that is enough to conclude the set is not closed.

myininaya (myininaya):

Think: What happens when you take a number and minus from it a greater number? like what is 5-7 or 0-1

OpenStudy (anonymous):

:/

myininaya (myininaya):

you can't tell me what 5-7 is?

OpenStudy (anonymous):

2